### v1

# DAMO: Data- and Model-aware Alignment of Multi-modal LLMs

Abstract
Direct Preference Optimization (DPO) has shown effectiveness in aligning multi-modal large language models (MLLM) with human preferences. However, existing methods exhibit an imbalanced responsiveness to the data of varying hardness, tending to overfit on the easy-to-distinguish data while underfitting on the hard-to-distinguish data. In this paper, we propose Data- and Model-aware DPO (DAMO) to dynamically adjust the optimization process from two key aspects: (1) a data-aware strategy that incorporates data hardness, and (2) a model-aware strategy that integrates real-time model responses. By combining the two strategies, DAMO enables the model to effectively adapt to data with varying levels of hardness. Extensive experiments on five benchmarks demonstrate that DAMO not only significantly enhances the trustworthiness, but also improves the effectiveness over general tasks. For instance, on the Object HalBench, our DAMO-7B reduces response-level and mentioned-level hallucination by 90.0% and 95.3%, respectively, surpassing the performance of GPT-4V. Code is available at: https://github.com/injadlu/DAMO
1 Introduction
Figure 1: (1) Preference data (Prompt, Image, Preferred response $y_{w}$, Rejected response $y_{l}$) with different hardness: “easy-to-distinguish” data denotes a large Image-Text sim score gap between $y_{l}$ and $y_{w}$; “hard-to-distinguish” data indicates a low score gap between $y_{l}$ and $y_{w}$. (2) Implicit reward across the optimization stage: the reward gap for “easy-to-distinguish” data enhances significantly during optimization, while for “hard-to-distinguish” data, the gap remains low.
Recent advances in Multimodal Large Language Models (MLLMs) have demonstrated remarkable visual understanding capabilities on the basis of large language models (Liu et al., 2024b; Chen et al., 2024; Bai et al., 2023). However, despite their effectiveness, the hallucination issue — generating outputs inconsistent with the image content and human preference — limits their broader applicability (Leng et al., 2024; Liu et al., 2023; Yu et al., 2024b). To address this, direct preference optimization (DPO) (Rafailov et al., 2024) has been adapted into MLLM alignment (Yu et al., 2024c; Wang et al., 2024; Ouali et al., 2025), achieving encouraging performance with moderate computational costs.
DPO methods (Sun et al., 2024; Yu et al., 2024c) collect preference data consisting of an image, a prompt, and two responses ($y_{w},y_{l}$). The preferred response $y_{w}$ is better aligned with the visual content (large Image-Text similarity score), while the rejected response $y_{l}$ contains more hallucinated content (small Image-Text similarity score). They prioritize preferred responses $y_{w}$ over the rejected ones $y_{l}$ using DPO with a hyperparameter $\beta$, which balances retaining the reference model $\pi_{\text{ref}}$ and incorporating new preferences into the updated model $\pi_{\bm{\theta}}$ (Gu et al., 2024; Xie et al., 2024).
However, our analysis reveals that current methods exhibit imbalanced responsiveness in handling the data with varying hardness during the optimization process, resulting in suboptimal performance. As illustrated in Figure 1, for “easy-to-distinguish” data (large Image-Text similarity score gap between preferred $y_{w}$ and rejected $y_{l}$), the reward gap amplifies during training, indicating stronger alignment. Conversely, for “hard-to-distinguish” data (small Image-Text similarity score gap), the reward gap stagnates, suggesting limited capability to distinguish $y_{w}$ from $y_{l}$. This implies that current methods, which employ optimization strategies using a static $\beta$ across data with varying hardness (Sun et al., 2024; Yu et al., 2024c), could fail to capture the learning dynamics inherent in multimodal preference data.
To address this imbalanced responsiveness issue, we propose Data- and Model-aware direct preference optimization (DAMO), which dynamically adapts $\beta$ to both data hardness and model’s responsiveness, enabling adaptively adjust model’s learning behavior based on the inherent preference dynamics and model’s real-time responses. Specifically, we propose two novel mechanisms:
Data-aware Preference Optimization. (Section 3.1): We quantify data hardness via CLIP-based image-text similarity scores (Radford et al., 2021), decomposing responses into sub-sentences for granularity-aware estimation. Then we normalize and transform the scores into probabilities to enable effective hardness estimation. By dynamically scaling $\beta$ inversely with hardness, we enforce stronger regularization on “easy-to-distinguish” samples (large $\beta$) while relaxing constraints for “hard-to-distinguish” ones (small $\beta$), preventing overfitting and underfitting, respectively.
Model-aware Preference Optimization. (Section 3.2): We estimate the model’s responsiveness through the reward gaps between the preferred and rejected responses. Moreover, to improve the robustness of the estimation, we normalize the reward gaps and filter out outliers. Similarly to Section 3.1, we incorporate the estimated responsiveness into the optimization process by dynamically modifying $\beta$ according to it, enabling the model to effectively optimize according to its current responsiveness.
By combining these strategies via element-wise multiplication, DAMO enables real-time adaptation to both data hardness and model responsiveness, demonstrating strong alignment performance across various evaluation benchmarks. Our contributions are summarized as follows:
  • We introduce a data-aware preference optimization strategy to integrate the data hardness into the preference optimization process, enabling the model to adaptively optimize based on the data hardness.
  • We propose a model-aware preference optimization strategy to incorporate the real-time model response into the preference optimization process, facilitating the model to effectively optimize based on its current responsiveness.
  • By combining the data- and model-aware strategies, we adaptively refine the optimization process, achieving significant alignment performance, as demonstrated by extensive empirical evaluations.
Figure 2: Overview of our data-aware preference optimization. For each preference instance: (1) We first break the preferred and rejected response into sub-sentences by prompting a large language model (LLM); (2) Next, we estimate the similarity scores between each sub-sentence and the given image using the CLIP classifier, and then calculate the differences between the preferred and rejected response as the hardness of the data; (3) Finally, we incorporate the estimated hardness into the preference optimization process by modifying $\beta$ in Equ (3), allowing the model to adjust based on the data hardness.
2 Preliminary
In this section, we briefly review the MLLM preference learning procedure, which starts by sampling pairwise preference data with a supervised fine-turned (SFT) model, and then optimizes on such preference data. Specifically, we categorize this process into the following aspects:
Supervised Fine-Tuning (SFT). Preference learning of an MLLM $\bm{\pi}$ begins with an SFT model $\bm{\pi}_{\text{SFT}}$. Concretely, the SFT process fine-tunes the pre-trained MLLM model with millions of multi-modal question-answer pairs to align LLM with multi-modal tasks. After this process, we construct preference data by sampling pair-wise preference responses from $\bm{\pi}_{\mathrm{SFT}}$, formalized as $(y_{w},y_{l})\sim\bm{\pi}_{\mathrm{SFT}}(y|x,\mathcal{I})$, where $(\mathcal{I}$ denotes the image and $x$ is the prompt question. Meanwhile, $(y_{w},y_{l})$ are labeled as preferred and less preferred responses by humans, formalized as $(y_{w}\succ y_{l}|\mathcal{I},x)$.
RLHF with Reward Models. Given pair-wise preference data $(y_{w},y_{l})\sim\bm{\pi}_{\mathrm{SFT}}(y|x,\mathcal{I})$, the preference learning process can be described in 2 stages: reward modeling and preference optimization. Specifically, the reward model $r_{\bm{\theta}}(y|\mathcal{I},x)$ is defined to rank the model responses by learning to distinguish $y_{w}$ from $y_{l}$, and the preference optimization aims to distill the preference knowledge into MLLM. To learn a reward model, pioneering work (Ouyang et al., 2022) employs the Bradley-Terry model (Bradley & Terry, 1952) to model the pair-wise preference distribution as:
[…]
Thus, the learning process can be achieved by minimizing the negative log-likelihood $-\mathrm{logP}(y_{w}\succ y_{l}|\mathcal{I},x)$ over the preference data with the parametrized reward model $r_{\bm{\phi}}(y_{w}|\mathcal{I},x)$ initialized as $\bm{\pi}_{\mathrm{SFT}}$ with a simple linear layer to produce reward prediction. With the well-optimized reward model $r_{\phi}^{*}(y|\mathcal{I},x)$, prior work (Ouyang et al., 2022) proposes to employ policy optimization algorithms in RL such as PPO (Schulman et al., 2017) to maximize the learned reward with KL-penalty, which can be formalized as:
$\underset{\bm{\pi}_{\theta}}{\text{max}}$ $\ \mathbf{E}_{(\mathcal{I},x)\sim\mathcal{D},y\sim\bm{\pi}_{\theta}(\cdot|\mathcal{I},x)}[r_{\phi}^{*}(y|\mathcal{I},x)]$ (2)
$-\beta\mathbb{D}_{\mathbf{KL}}[\bm{\pi}_{\theta}(y|\mathcal{I},x)||\bm{\pi}_{\text{ref}}(y|\mathcal{I},x)],$
where the fixed reference model $\bm{\pi}_{\text{ref}}$ is parameterized as $\bm{\pi}_{\text{SFT}}$, and the hyper-parameter $\beta$ controls the deviation of $\bm{\pi}_{\theta}$ from $\bm{\pi}_{\text{ref}}$ during the optimization process.
Direct Preference Optimization (DPO). To relieve the high computational complexity of reward training in RLHF, DPO (Rafailov et al., 2024) is proposed, which provides a simple way to directly optimize $\bm{\pi}_{\theta}$ with the pair-wise preference data, without parametrized reward model. Specifically, the DPO loss can be described as:
$\mathcal{L}_{\mathrm{dpo}}=-\bm{\mathrm{E}}_{(\mathcal{I},x,y_{w},y_{l})}[{\log\sigma}($ $\beta\log\frac{{\pi}_{\bm{\theta}}(y_{w}|\mathcal{I},x)}{{\pi}_{\mathrm{ref}}(y_{w}|\mathcal{I},x)}$ (3)
-- $\beta\log\frac{{\pi}_{\bm{\theta}}(y_{l}|\mathcal{I},x)}{{\pi}_{\mathrm{ref}}(y_{l}|\mathcal{I},x)})].$
3 Approach
In this section, we describe our DAMO in detail. Specifically, we first illustrate our data-aware preference optimization, then we describe our model-aware preference optimization, and finally, we show our combination strategies for robust preference optimization. Our approach algorithm is listed in Algorithm. 1.
3.1 Data-aware Preference Optimization
An overview of our data-aware preference optimization is shown in Figure 2. Given a preference instance from the dataset $\mathcal{D}$ as $\{(\mathcal{I},x,y_{w},y_{l})\}\sim\mathcal{D}$, where $\mathcal{I}$, $x$, $y_{w}$, $y_{l}$ denotes the image, question, preferred response, and rejected response, respectively, it firstly splits the responses into simple and self-contained sub-sentences. Next, it calculates the image-text similarity scores between the sub-sentences and the image by the CLIP classifier. Then, it combines the scores of each response and compares the difference between the preferred and rejected responses as the data hardness. This hardness is embedded into the preference optimization process by modifying the $\beta$ in Equ (3). The following are detailed descriptions.
We employ the CLIP classifier $\mathbf{\Gamma}_{\mathrm{CLIP}}$ (Radford et al., 2021), to calculate similarity scores. For each preference instance, we aim to effectively capture the similarity between the responses $(y_{w},y_{l})$ with the given image $\mathcal{I}$, while alleviating the 77 token length constraints in CLIP. To achieve this, we decompose the complex responses, which contain various objects and relations, into simple and self-contained sub-sentences. Concretely, we prompt the open-source large language model, such as LLaMA-3 (Dubey et al., 2024), to split $(y_{w},y_{l})$ into sub-sentences $\mathbf{S}_{w}=\{\mathbf{S}_{w,j}|j=1,2,\dots p\}$ and $\mathbf{S}_{l}=\{\mathbf{S}_{l,k}|k=1,2,\dots q\}$, where $p$ and $q$ denotes the number of sub-sentences for $\mathbf{S}_{w}$ and $\mathbf{S}_{l}$.
Subsequently, we employ the CLIP classifier $\mathbf{\Gamma}_{\mathrm{CLIP}}$ to calculate the similarity score between the given image $\mathcal{I}$ and the sub-sentences $\mathbf{S}_{w},\mathbf{S}_{l}$ as:
$\mathbf{C}_{w}$ $=\left[\mathbf{\Gamma}_{\mathrm{CLIP}}(\mathcal{I},\mathbf{S}_{w,j})\right]_{j=1}^{p},$ (4)
$\mathbf{C}_{l}$ $=\left[\mathbf{\Gamma}_{\mathrm{CLIP}}(\mathcal{I},\mathbf{S}_{l,k})\right]_{k=1}^{q},$
where $\mathbf{C}_{w}\in\mathbb{R}^{p}$ and $\mathbf{C}_{l}\in\mathbb{R}^{q}$ represents the corresponding similarity scores of $\mathbf{S}_{w}$ and $\mathbf{S}_{l}$, respectively. To effectively quantify the difference between preferred response $\mathbf{C}_{w}$ and rejected response $\mathbf{C}_{l}$ for each instance, we normalize the corresponding score by the softmax probabilities as :
$\left[\begin{array}[]{c}\mathbf{P}_{w}\\ \mathbf{P}_{l}\end{array}\right]=\mathbf{Softmax}\left(\left[\begin{array}[]{c}\mathbf{C}_{w}\\ \mathbf{C}_{l}\end{array}\right]\right),$ (5)
$\mathbf{P}_{w}\in\mathbb{R}^{p}$ and $\mathbf{P}_{l}\in\mathbb{R}^{q}$ represents the probabilites. The difference between the preferred and rejected probabilities demonstrates the data hardness. A large difference implies that the preference data is “easy-to-distinguish”, where the rejected response includes more elements that are not present in the image, conversely, a small difference suggests that the preference data is “hard-to-distinguish”, and the rejected response exhibits minimal hallucination. Then we define the hardness based on the probabilities difference as:
$\delta=$ $\sum_{j=1}^{p}\mathbf{P}_{w,j}-\sum_{k=1}^{q}\mathbf{P}_{l,k},$ (6)
$\alpha_{\mathrm{D}}=\sigma(\delta)/\sigma(\bar{\delta}),$ (7)
where $\alpha_{\mathrm{D}}$ denotes the data hardness, $\delta$ measures the difference between $\mathbf{P}_{w}$ and $\mathbf{P}_{l}$, and $\bar{\delta}$ denotes the mean difference across the dataset. The Sigmoid function $\sigma(\cdot)$ is employed to transform the response divergence $\mathbf{P}_{w}$ and the mean $\bar{\delta}$ into the range $(0,1)$ for convenient comparison.
Finally, we adapt $\beta$ of Equ (3) to incorporate the hardness into the optimization procedure, and each preference instance corresponding to a specific $\beta$ as:
$\beta_{\mathrm{D}}=\beta\cdot\alpha_{\mathrm{D}}.$ (8)
This adjustment allows the model to optimize based on the data hardness, further enhancing its adaptability to the data.
Figure 3: Overview of our model-aware preference optimization. Given $N$ preference instances: (1) we first calculate the reward gap of each instance using the implicit reward model; (2) To ensure stable modeling, we filter out the outliers (i.e. the instance with excessively high or low gaps) and then estimate the average gap; (3) To enable the model to be aware of its current responsiveness, we integrate such estimation into the preference optimization process by modifying $\beta$ in Equ (3).
3.2 Model-aware Preference Optimization
An overview of our model-aware preference optimization is shown in Figure 3. Given a batch of preference instances from dataset $\mathcal{D}$, as =$\mathcal{B}=\{(\mathcal{I}_{i},x_{i},y_{w,i},y_{l,i})|i=1,2,\dots,N\}\sim\mathcal{D}$, it firstly calculates the reward gaps between the preferred $y_{w,i}$ and rejected $y_{l,i}$, and then filters out the outliers (i.e. the instance with excessively high or low gaps) for stable estimation. Such estimations are embedded into the preference optimization process by integrating into $\beta$ in Equ (3), enabling the model to be aware of its current responsiveness. Details of model-aware preference optimization are as follows.
We employ current implicit reward gaps between the preferred and rejected responses of the given $\mathcal{B}$ instances to measure the current model’s responsiveness. Specifically, the reward gap $\mathcal{R}$ for the $i$-th instance in $\mathcal{B}$ is formalized as:
$\mathcal{R}_{i}=\left[\beta\log\frac{{\pi}_{\bm{\theta}}(y_{w,i}|\mathcal{I}_{i},x_{i})}{{\pi}_{\text{ref}}(y_{w,i}|\mathcal{I}_{i},x_{i})}-\beta\log\frac{{\pi}_{\bm{\theta}}(y_{l,i}|\mathcal{I}_{i},x_{i})}{{\pi}_{\text{ref}}(y_{l,i}|\mathcal{I}_{i},x_{i})}\right],$
where ${\pi}_{\bm{\theta}}$ and ${\pi}_{\text{ref}}$ represent the optimizing model and reference model, respectively. We then normalize the reward gaps using the estimated mean as follows:
$\bar{{\mathcal{R}}_{i}}=\mathcal{R}_{i}/\bar{\mathcal{R}},$ (10)
where $\bar{\mathcal{R}}$ represents the estimated average reward gap, and $\bar{{\mathcal{R}}_{i}}$ is the normalized one for the $i$-th instance.
However, the estimation remains sensitive to outliers despite normalization, especially in the full fine-tuning scenario, where the batch size is relatively small. To mitigate this issue, we filter out instances with exceptionally high or low gaps using a mask vector $\mathcal{M}\in\mathbb{R}^{N}$, defined as:
$\bm{M}_{i}=\left\{\begin{aligned} &1,&(\bar{\mathcal{R}}_{i}-\bar{\mathcal{R}})^{2}\leq\tau,\\ &0,&(\bar{\mathcal{R}}_{i}-\bar{\mathcal{R}})^{2}>\tau,\end{aligned}\right.$ (11)
where $(\bar{\mathcal{R}_{i}}-\bar{\mathcal{R}})^{2}$ implies the squared distances from the mean, and $\tau$ represents the sorted $K$-th distance. With the filtering, current responsiveness of ${\pi}_{\bm{\theta}}$ can be formalized as:
$\bar{\mathcal{R}}_{\mathcal{B}}$ $=\frac{1}{N-K}\sum_{i=1}^{N}\mathcal{M}_{i}\times\bar{\mathcal{R}}_{i},$ (12)
$\alpha_{\text{M}}$ $=\sigma(\bar{\mathcal{R}}_{\mathcal{B}})/\sigma(\bar{\mathcal{R}}).$ (13)
$\sigma(\cdot)$ is the Sigmoid function, which transforms both the filtered gaps $\bar{\mathcal{R}}_{\mathcal{B}}$ and the estimated mean $\bar{\mathcal{R}}$ into the range $(0,1)$ for convenient comparison, and $\alpha_{M}$ refers to the estimated model responsiveness.
We then integrate $\alpha_{M}$ into the optimization procedure by modifying $\beta$ of Equ (3) as:
$\beta_{\text{M}}=\beta\cdot\alpha_{\text{M}}.$ (14)
By utilizing $\beta_{M}$ to optimize $\pi_{\theta}$ with the filtered batch $\mathcal{B}\cdot\mathbf{M}$, the model can effectively adapt to its current responsiveness to the preference data.
Finally, we update the estimated mean $\bar{\mathcal{R}}$ using a moving average after optimization over batch $B$ as follows:
¯$\bar{\mathcal{R}}\bm{\leftarrow}\gamma\cdot\bar{\mathcal{R}}+(1-\gamma)\cdot\bar{\mathcal{R}}_{\mathcal{B}},$ (15)
the momentum $\gamma$ is set to $0.9$, and $\bar{\mathcal{H}}$ is initialized to 00.
3.3 Combining for Preference Optimization
In this section, we present the combination of both data- and model-aware strategies for robust preference optimization. Specifically, given a batch of preference instances =$\mathcal{B}=\{(\mathcal{I}_{i},x_{i},y_{w,i},y_{l,i})|i=1,2,\dots,N\}\sim\mathcal{D}$, where $\mathcal{I}_{i}$, $x_{i}$, $y_{w,i}$, $y_{l,i}$ denotes the image, question, preferred response, and rejected response, respectively, we first compute data hardness offline based on our data-aware preference optimization, and then obtain the corresponding instance-wise hardness by Equ (7) as $\alpha_{\mathrm{D}}^{\mathcal{B}}\in\mathbb{R}^{N}$. Next, we calculate the reward gaps using Equ 13 as $\alpha_{\text{M}}$. To facilitate the optimization process to be both data- and model-aware, we propose an element-wise combination strategy. Specifically, we combine $\alpha_{\mathrm{D}}^{\mathcal{B}}$ and $\alpha_{\text{M}}$ as:
$\alpha=\alpha_{\mathrm{D}}^{\mathcal{B}}\cdot\alpha_{\text{M}},$ (16)
where $\alpha\in\mathbb{R}^{N}$ represents the combined factor. Subsequently, we adjust $\beta$ of Equ (3) to incorporate both components into optimization procedure as:
$\beta_{\text{C}}=\beta\cdot\alpha,$ (17)
where $\beta_{\text{C}}\in\mathbb{R}^{N}$, and each preference instance in $\mathcal{B}$ corresponds to a specific $\beta$. Finally, our combination strategy can be achieved by employing the $\beta_{\text{C}}$ to optimize $\pi_{\bm{\theta}}$ with the filtered batch $\mathcal{B}\cdot\bm{M}$, where $\bm{M}$ is obtained by Equ (11). Thus, the optimization process can become more adaptive, allowing the model to refine its preferences based on both pre-computed data hardness and real-time model responsiveness, further enhancing the robustness.
Algorithm 1 Algorithm of DAMO.
  Input: Preference dataset $\mathcal{D}$, hyper-parameter $\beta$, SFT model $\pi_{\text{SFT}}$, CLIP classifier $\Gamma_{\text{CLIP}}$.
  Output: The optimized model $\pi_{\bm{\theta}}$.
  Initialize model $\pi_{\bm{\theta}}$ and reference model $\pi_{\text{ref}}$ as $\pi_{\text{SFT}}$.
  for $\{(\mathcal{I},x,y_{w},y_{l})\}$ in $\mathcal{D}$ do
     $\mathbf{S}_{w}\leftarrow\text{LLM}\{y_{w}\}$, $\mathbf{S}_{l}\leftarrow\text{LLM}\{y_{l}\}$;
     obtains $\delta$ with $\{\mathcal{I},\mathbf{S}_{w}\}$, $\{\mathcal{I},\mathbf{S}_{l}\}$; $\triangleright$ Equ (4) $\to$ (6);
     $\alpha_{\mathrm{D}}\leftarrow\sigma(\delta)/\sigma(\bar{\delta})$; $\triangleright$ Equ (7);
  end for
  repeat
     for =$\mathcal{B}=\{(\mathcal{I}_{i},x_{i},y_{w,i},y_{l,i})\}_{i=1}^{N}\sim\mathcal{D}$ do
        obtain $\mathcal{R}_{i}$ with $y_{w,i}$ and $y_{l,i}$; $\triangleright$ Equ (9);
        obtain $\bar{\mathcal{R}}_{\mathcal{B}}$ with $\mathcal{R}_{i}$; $\triangleright$ Equ (10) $\to$ (12);
        $\alpha_{M}\leftarrow\sigma(\bar{\mathcal{R}}_{\mathcal{B}})/\sigma(\bar{\mathcal{R}})$; $\triangleright$ Equ (13);
        $\alpha\leftarrow\alpha_{\mathrm{D}}^{\mathcal{B}}\cdot\alpha_{M}$, where $\alpha_{\mathrm{D}}^{\mathcal{B}}=\{{\alpha_{\mathrm{D,i}}}\}_{i=1}^{N}$; $\triangleright$ Equ (16);
        $\beta_{\mathrm{C}}\leftarrow\beta\cdot\alpha$; $\triangleright$ Equ (17);
        Compute loss w.r.t. $\beta_{\mathrm{C}},\pi_{\bm{\theta}}$; $\triangleright$ Equ (3);
        Compute the gradient and update the model $\pi_{\bm{\theta}}$.
        ¯$\bar{\mathcal{R}}\bm{\leftarrow}\gamma\cdot\bar{\mathcal{R}}+(1-\gamma)\cdot\bar{\mathcal{R}}_{\mathcal{B}}$; $\triangleright$ Equ (15);
     end for
  until The optimization is converged.
4 Experiment
In this section, we elaborate on the effectiveness of our Data- and Model-aware Direct Preference Optimization (DAMO). Specifically, we first introduce the details of our experimental settings. Next, we illustrate the ablation studies, and finally, we compare the results with the state-of-the-art methods over various benchmarks.
4.1 Experimental Settings
In this section, we describe the experimental settings.
Backbone: We employ the LLaVA-1.5 7B and 13B for performance comparison (Liu et al., 2024b).
Dataset: Our focus is not on the preference data construction, thus we directly utilize the released dataset by (Yu et al., 2024c), which contains 22k preference data totally.
Baselines. In this work, we compare against state-of-the-art baselines across various categories:
(1) Hallucination-specific baselines. In this category, we mainly compare with VCD (Leng et al., 2024), Less-is-more (Yue et al., 2024), OPERA (Huang et al., 2024), and CCA-LLaVA (Xing et al., 2024).
(2) Preference Optimization-based baselines. In this category, we mainly compare with HA-DPO (Zhao et al., 2023), POVID (Zhou et al., 2024a), LLaVA-RLHF (Sun et al., 2024), RLHF-V (Yu et al., 2024b), RLAIF-V (Yu et al., 2024c), AMP-MEG (Zhang et al., 2024a), CSR (Zhou et al., 2024b), V-DPO (Xie et al., 2024), and TPO (Gu et al., 2024).
(3) Commercial baseline. We include GPT-4V as a strong reference to evaluate the performance gap between the open-source and commercial models.
Benchmarks. We conduct experiments on five benchmarks, including three hallucination benchmarks reflecting trustworthiness, and two general benchmarks:
(1) Object HalBench is employed to evaluate object hallucination by detailed descriptions of the image content. We report the response-level and mentioned-level non-hallucination rates to evaluate its capability to reduce hallucination (Rohrbach et al., 2018).
(2) AMBER is a multi-dimensional hallucination benchmark, which contains more than 15k samples. We report the Accuracy and F1 metric by its discriminative component (Wang et al., 2023).
(3) MMHal-Bench assesses response-level hallucination rate and informativeness by GPT-4 compare model outputs with human responses and object labels (Sun et al., 2024).
(4) LLaVA Bench consists of 24 images and 60 questions including conversation, detailed description, and complex reasoning ability (Liu et al., 2024b).
(5) MM-Vet is designed to evaluate six integrated competencies, including OCR, recognition, knowledge, language generation, spatial awareness, and math (Yu et al., 2023).
Implementation Details. For both LLaVA-1.5 7B and 13B models, we employ full parameter-tuning over the preference dataset with four epochs. Specifically, for reproducibility, we adopt the same hyperparameters as provided in the official LLaVA GitHub repository  [^1] https://github.com/haotian-liu/LLaVA. The batch size $N$ is set to $16$, the selected size $K$ is set to $12$, and the penalty hyperparameter $\beta$ is set to $0.1$ by following (Rafailov et al., 2024; Yu et al., 2024c). All experiments are conducted with four A100 80GB GPUs, and four epochs of fine-tuning cost seven hours for both backbones.
4.2 Ablation Studies
In this section, we evaluate the effects of different components of our DAMO. To this end, we utilize the LLaVA-1.5 7B model as the backbone. For clear illustration, we report both the response and the mentioned-level non-hallucination rate on the Object Hallucination benchmark, where the non-hallucination rate is defined as $100\%-$ hallucination rate. The following are detailed illustrations.
Influences of different components. The performance of various components of DAMO is reported in Table 1, where “DPO” refers to Direct Preference Optimization (Rafailov et al., 2024), “MDPO” represents our Model-aware Preference Optimization, “D2PO” denotes Data-aware Preference Optimization, and “DAMO” corresponds to our combined strategy. The experimental results demonstrate: (1) All strategies significantly outperform the baseline method (DPO), especially our final “DAMO”, achieving more than $10\%$ response level performance gains, highlighting the effectiveness of our method; (2) Compared with “MDPO”, the performance gain of “D2PO” is relatively modest, suggesting that the quality of the preference data is already high, which further validates the efficacy of the preference data construction strategy (Yu et al., 2024c).
Table 1: Experimental results of different components of DAMO.
Method Object HalBench
Response ($\uparrow$) Mention ($\uparrow$)
LLaVA-1.5-7B 47.75 73.08
+DPO 78.29 89.48
+D2PO        82.54
+4.25
       90.64
+1.16
+MDPO        88.00
+9.71
       93.74
+4.26
+DAMO         90.87 +12.58        95.33 +5.85
Influences of probability transformation in data-aware preference optimization. Table 2 summarizes the performance of different response inconsistency estimation strategies, where “CLIP Scores” denotes that we directly estimate $\delta$ based on $\mathbf{C}_{w}$ and $\mathbf{C}_{l}$, with $\delta=\sum_{j=1}^{p}\mathbf{C}_{w,j}/\sum_{k=1}^{q}\mathbf{C}_{l,k}$, and “CLIP Probs” represents our strategy, which transforms $\mathbf{C}_{w}$ and $\mathbf{C}_{l}$ into probabilities. The results indicate that: firstly, both strategies improve the performance, underscoring the effectiveness of integrating data inconsistency into the optimization process, which allows the model to better handle varying levels of data hardness, thereby improving overall robustness; furthermore, transforming $\mathbf{C}{w}$ and $\mathbf{C}_{l}$ into probabilities yields a larger performance gain, as probabilities smooth the estimation of $\delta$, mitigating noise from large gaps between $\mathbf{C}{w}$ and $\mathbf{C}_{l}$, and preventing the influences by the outliers.
Table 2: Experimental results of different preference inconsistency construction strategies.
Method Object HalBench
Response ($\uparrow$) Mention ($\uparrow$)
DPO 78.29 89.48
+CLIP Scores        80.65
+2.36
       89.62
+0.14
+CLIP Probs        82.54 +4.25        90.64 +1.16
Effects of the data filtering in model-aware preference optimization. Table 3 presents the performance of different data filtering strategies, where “No Filter” denotes that we directly utilize the mean gaps to estimate the model state without filtering, “Bottom” shows that we remove the $N-K$ samples with the largest distances in the batch, “Top” is filtering the $N-K$ samples with the smallest distances, “Bottom & Top” refers to our filtering strategy, which filters both extremes based on the squared distances. Specifically, we can observe that: firstly, filtering solely from the bottom or top leads to performance degradation, indicating that such data introduces bias in estimating the model state. Moreover, filtering only the bottom samples results in significant performance drops due to overfitting on the top-ranked data, which misguides the estimation of model responsiveness to focus excessively on potentially less representative instances. Furthermore, filtering both bottom and top samples yields performance improvements, demonstrating the effectiveness of our proposed strategy, as it balances the influence of extreme data points.
Table 3: Experimental results of different filtering strategies.
Method Object HalBench
Response ($\uparrow$) Mention ($\uparrow$)
No Filter 86.66 92.62
Bottom        76.36
-10.30
      87.20
-5.42
Top       86.34
-0.32
      92.48
-0.14
Bottom & Top        88.00 +1.34        93.74 +1.12
Table 4: Performance comparisons with state-of-the-art methods on different benchmarks. We report non-hallucination rates in different levels including response level (Non-Rsp.) and mentioned-level (Non-Men.) for Object HalBench (Rohrbach et al., 2018). Hall. refers to the Hallucination Rate for MMHal Bench (Sun et al., 2024). The best results of all methods are indicated in bold, and the second best results are underlined. The compared results are sourced from (Yu et al., 2024c; Gu et al., 2024), and the reported results of LLaVA-1.5, DPO, and DAMO are evaluated using GPT-4-turbo-2024-04-09.
Method Size Object AMBER MMHal- LLaVA MM-Vet($\uparrow$)
HalBench Bench Bench($\uparrow$)
Non-Rsp.($\uparrow$) Non-Men.($\uparrow$) Acc($\uparrow$) F1($\uparrow$) Scores($\uparrow$) Hall.($\downarrow$)
Method (Hallucination-specific)
VCD
(CVPR’24)
7B 51.2 75.7 71.8 74.9 2.12 54.2 61.6 -
Less-is-more
(ACL’24)
7B 59.7 82.2 72.4 75.8 2.33 50.0 - -
OPERA
(CVPR’24)
7B 54.9 77.7 75.2 78.3 2.15 54.2 61.3 -
CCA-LLaVA
(NIPS’24)
7B 53.3 76.2 77.7 81.9 1.92 61.5 64.3 -
Method (Preference optimization)
HA-DPO
(arXiv’23)
7B 60.1 80.1 75.2 79.9 1.98 60.4 - -
POVID
(arXiv’24)
7B 51.9 75.6 82.9 87.4 2.08 56.2 68.2 31.7
RLHF-V
(CVPR’24)
7B - - 74.8 78.5 2.02 60.4 68.0 32.3
RLAIF-V
(arXiv’24)
7B 89.5 94.8 76.8 84.5 2.95 32.3 - -
CSR
(NIPS’24)
7B - - 73.2 76.1 2.05 60.4 68.9 31.0
V-DPO
(EMNLP’24)
7B - - - 81.6 2.16 56.0 - -
TPO
(arXiv’24)
7B - - 79.3 85.0 2.47 51.0 70.2 33.0
TPO
(arXiv’24)
13B - - 83.9 88.0 2.72 45.8 72.8 36.2
LLaVA-RLHF
(ACL’24)
13B 61.9 81.1 79.7 83.9 2.02 62.5 95.6 -
RLHF-V
(CVPR’24)
13B 87.8 92.5 72.6 75.0 2.45 51.0 76.7 38.5
AMP-MEG
(NIPS’24)
13B 68.3 79.4 79.5 84.6 3.08 36.5 - -
LLaVA-1.5 7B 47.8 71.2 73.9 77.7 1.95 63.5 62.3 31.6
+ DPO 7B 78.3 89.5 75.5 79.2 2.15 57.0 64.6 30.4
+ DAMO 7B 90.9 95.3 83.3 87.0 2.76 41.0 68.0 32.8
Improvements (%) +16.1% +6.5% +10.3% +9.8% +28.4% +28.1% +5.3% +7.9%
LLaVA-1.5 13B 50.0 76.4 71.2 73.0 2.36 56.0 66.1 36.1
+ DPO 13B 84.5 92.4 78.5 83.5 2.59 48.0 74.0 35.4
+ DAMO 13B 89.1 94.4 84.3 88.1 2.89 43.0 75.1 36.4
Improvements (%) +5.1 % +2.1 % + 5.8% + 4.6% +15.4% +10.4% +1.4 % +2.8 %
GPT-4V - 86.4 92.7 83.4 87.4 3.49 28.1 98.0 67.7
Effects of the combination strategy. Figure 4 and Figure  5 illustrate the performance of different combination strategies from the Response and Mention, respectively. “Multiplication” refers to combining both effects using a weighted sum strategy, where $\alpha=(1-\rho)\cdot\alpha_{M}+\rho\cdot\alpha_{\mathrm{D}}^{\mathcal{B}}$, with $\rho$ ranging from $0.0$ to $1.0$, where $\rho=0.0$ is the MDPO, and $\rho=1.0$ refers to the “D2PO”. We can observe that multiplication outperforms the weighted sum by a significant margin. This is because the weighted sum balances the two components without effectively merging their influences, whereas multiplication amplifies their contributions, allowing for more robust and superior performance gains.
Figure 4: Experimental results of the combination strategies with the response-level non-hallucination rates.
Figure 5: Experimental results of the combination strategies with the mentioned-level non-hallucination rates.
4.3 Comparison with state-of-the-art methods
In this subsection, we compare our method with state-of-the-art methods under three trustworthy benchmarks: Object HalBench, AMBER, and MMHal-Bench, and two general benchmarks: LLaVA-Bench and MM-Vet. We compare our method against baselines from various aspects, including hallucination-specific baselines, preference optimization-based baselines, and the commercial baseline GPT-4V. Our DAMO is adapted to both the LLaVA-1.5 7B and 13B models. The experimental results, along with results of vanilla DPO and the improvements achieved by DAMO over DPO, are listed in Table 4.
From the experimental results, we can observe that: (1) Compared to DPO, our DAMO achieves substantial performance improvements over all compared benchmarks. It’s noteworthy that on MMHal-Bench, it achieves more than 10% improvements for the 13B model, and more than 28% gain for 7B models, which is significant; (2) DAMO achieves new state-of-the-art over various benchmarks. Our DAMO-7B reduces the response-level and mentioned-level hallucination on the generative Object HalBench by 90.9% and 95.3%. Moreover, DAMO-13B achieves 84.3% Accuracy and 88.1 F1 score on the discriminative AMBER benchmark. These attained results surpass those of GPT-4V, fully demonstrating the effectiveness of DAMO.
5 Related Work
In this section, we overview the related background. Specifically, we first briefly summarize recent methods for aligning with human preferences, and then, we discuss existing hallucination mitigation methods for multi-modal large language models. Finally, we enumerate the differences between ours and related methods.
5.1 Alignment with Human Preference
To align the model outputs with human preferences, preference optimization methods have garnered significant attention (Ouyang et al., 2022; Rafailov et al., 2024; Meng et al., 2024). Concretely, RLHF ((Ouyang et al., 2022)) first introduces alignment through reward modeling, which trains a parametric reward model on preference data and subsequently optimizes the preference model using PPO (Schulman et al., 2017). However, obtaining an effective reward model is challenging. To address this, DPO (Rafailov et al., 2024) is proposed to simplify the reward modeling process with an implicit reward function, allowing for direct optimization of preference data. Subsequent works like (Meng et al., 2024) and (Ethayarajh et al., 2024) further simplify this process by removing the reference model and introducing binary feedback, respectively.
Current alignment works in MLLMs focus on two aspects: (1) Collecting high-quality preference data. For example, existing works (Yu et al., 2024b, c; Li et al., 2024) construct high-quality preference data to advance the research of MLLM alignment. For example, the works in (Yu et al., 2024b; Sun et al., 2024) introduce human-based preference construction strategies, RLAIF-V (Yu et al., 2024c) utilizes the open-source models (e.g. LLaVA-NeXT (Liu et al., 2024a)) to construct preference data, and the works in (Li et al., 2024) and (Zhang et al., 2024b) employ closed-source models like GPT-4V for preference annotation. (2) Emphasizing the focus on visual detail. For instance, the works in (Xie et al., 2024) and (Wang et al., 2024) construct rejected samples by destroying the image, the works like TPO (Gu et al., 2024) and FiSAO (Cui et al., 2024) identify and emphasize key tokens from the preference responses to attend to visual detail.
5.2 Hallucination mitigation in MLLMs
Hallucination, As a key indicator of trustworthiness, refers to that the MLLM outputs are not aligned with the image content (Liu et al., 2024c). Beyond the preference optimization-based methods, current works to mitigate hallucination can be summarized into the following three aspects: (1) Data filtering. Works in LRV-Instruction (Liu et al., 2023) and Hallucidoctor (Yu et al., 2024a) identify that noises within the instruction tuning data serve as a cause of hallucination, and introduce fine-grained data cleaning and curation strategies to mitigate hallucination. (2) Enhanced visual representation. Works like (Jain et al., 2024; Tong et al., 2024) suggest that insufficient visual cues are the fundamental cause of hallucination, and they incorporate more intricate visual features to enrich the visual representations. (3) Inference-time enhancement. Works like VCD (Leng et al., 2024) and MARINE (Zhao et al., 2024) introduce visual contrastive decoding mechanisms to help enhance the model’s focus on visual details by contrastively sampling from original and visually distorted distribution.
Compared with existing methods, which focus on data curation and introduce fine-grained regularizations, our proposed method aims to optimize adaptively based on the specific model responsiveness and the hardness of the data. We seek to design an improved optimization strategy that can effectively leverage the current model and the preference data, enabling the model to adaptively respond to the preference data, and thereby enhancing the alignment performance.
6 Conclusion
In this paper, to address the imbalanced responsiveness to data with varying hardness, we propose Data- and Model-aware Direct Preference Optimization (DAMO) to adaptively adjust the model’s learning behavior from two aspects: (1) a data-aware strategy is proposed to dynamically adapt by incorporating the data hardness; and (2) a model-aware strategy is introduced to adaptively adjust by integrating the model’s real-time responsiveness. Extensive experiments on five benchmarks from distinct aspects demonstrate the effectiveness of our method.
Limitation and Future Work. Our focus is on MLLMs handling images and text; future work should explore scalability to broader applications, including video and audio processing, as well as commercial multi-modal large language models, to fully demonstrate its practical impact.
Impact Statement
This paper aims to enhance the robustness of multi-modal large language model alignment, thereby advancing the field of visual understanding. By tackling challenges like hallucination, we strive to improve the reliability of AI systems in alignment with human preferences. We hope that our research will contribute to facilitating the development of AI systems that are both effective and reliable, ultimately delivering meaningful societal benefits.
Appendix A Appendix
A.1 Qualitative analysis
In this section, we provide qualitative analysis between our Data- and Model-aware preference optimization (DAMO) and the DPO method (Rafailov et al., 2024). The case studies are shown in Figure 6 and Figure 7, and we also include the evaluations of GPT-4. From the case studies, we can observe two key points: (1) while both methods generate correct outputs, DPO tends to be more general, whereas DAMO offers richer, more detailed descriptions. This can be attributed to the integration of data hardness, quantified by the CLIP score between the Image and responses, into the optimization process, enabling DAMO to capture and describe more nuanced visual details; and (2) as evidenced by the GPT-4 evaluation, DAMO generates more vivid and immersive visual descriptions, this can be attributed to the enhancement of the responsiveness to the visual content, resulting in a more expressive and contextually aligned representation of the scene.
Figure 6: Qualitative Comparisons between DAMO and vanilla DPO. Data source: Object HalBench.
Figure 7: Qualitative Comparisons between DAMO and vanilla DPO. Data source: Object HalBench.